Subject: APO 3.5m users committee, Feb 2012 minutes

From: Suzanne Hawley

Submitted: Tue, 6 Mar 2012 14:45:43 -0800 (PST)

Message number: 1264 (previous: 1263, next: 1265 up: Index)

 		APO 3.5-m Users Committee Phonecon, 2/27/12

Attending:  Suzanne Hawley, Mark Klaene, Bill Ketzeback, Sean Moran, 
Jon Holtzman, John Bally, Michael Strauss, John Wilson, Matt Nelson, 
and Bruce Gillespie

Absent:  Remy Indebetouw, Scott Anderson, Al Harper


User feedback and comments from institutional representatives:

   o Chicago (Al Harper) - No report.

   o Washington (Scott Anderson) - Suzanne, for Scott, had nothing to report.

   o New Mexico State (Jon Holtzman) - Jon had nothing to report.

   o Princeton (Michael Strauss) - Michael said everything is fine.

   o Colorado (John Bally) - John had nothing to report.

   o Virginia (Remy Indebetouw) - No report.

   o Johns Hopkins (Sean Moran) - Sean had nothing to report.


Discussion of telescope/instruments report:

   Mark announced that the Wyoming NIIS instrument is on site, cooling, 
and appears to be ready for its trial runs on the telescope.  As mentioned 
before, if any ARC users have targets of interest to try, they should get 
in touch will Bill Ketzeback (

   The new B400 grating for DIS will likely be installed sometime in March. 
There have been vacuum problems with DIS blue camera, and its ion pump 
will be replaced to see if the problems go away.  SPIcam has developed 
some contamination on its detector, which seems to be going away slowly. 
The detector is being warmed slightly to accelerate this.  Users should 
consult the telescope and instrument news webpage at

for updates.  Michael asked if SPIcam is usable in its current state, and 
Bill said it should be, but users would be wise to take frequent calibrations. 
We will warm and pump the instrument when the weather gets warmer -- it's 
risky to do this in cold weather because of the possibility that the 
cryotiger will fail to restart properly.

   The pattern noise in TripleSpec data appears to have been essentially 
eliminated with the installation of a new power supply, but testing will 
continue.  The problem with the slit mechanism will be discussed separately; 
see below.

   The new echelle slitviewer camera is operational, and we are seeing
improvements in sensitivity.  The site engineering staff is checking on a 
report that slitviewer detector defects may interfere with target acquisition. 
The replacement grating for the echelle is in hand, and is planned for 
installation later in the spring.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                 1/17/11 through 02/22/12

0) Overview

Telescope continued to run well and with few instrument problems.  We continued exploring the scattered light performance out to 12'x12' FOV at NA2.  We also installed the new ECam.  The Wyoming NIIS wide-field NIR imager arrived on site and is being prepared for its inaugural observing run.

1) Telescope

The telescope has been operational with no issues of note.

2) Instruments

DIS replacement Blue 400 grating was tested at JHU and found to have slightly improved performance from the Newport data delivered with the instrument.  The grating is on site and will be installed in the instrument at the next convenient time.  We did have a short instance of vacuum level rising; a quick pump down allowed the instrument to continue observing.

Agile and SPIcam operated without issues.

TSpec pattern noise has been greatly improved with a new power supply and some tweaking by Ed.  Since the new power supply has been installed the pattern noise has not reappeared and we are tentatively calling this problem solved.  A problem with the slit mechanism did arise again and appears to be in the dewar.  Currently the slit is locked in the 1.1" position. see for up-to-date information.  (Will discuss slit issue in more detail later in agenda.)

Echelle is operational.  Plans for installation of the new grating are being developed.  Possible installation is May.  The new ECam was installed during Feb engineering time.  This required a lot of effort by Fritz, Bill, Joe, and Nick as well as the observing specialists to get it operational and ready for observing in such a short time and with minimal disruption to scheduled science.  The improvements in QE as well as other changes/improvements are documented in the Echelle User's Guide online documentation:

NIC-FPS has been operating in shared-risk observing status.  The image corruption issues are still present on occasion but have not been reported to be affecting science observations.

GFP software development by Jon Brinkmann et al. allowed for a moderately successful observing run this month.  Next step is refinement and TUI integration for the basic functionalities.


TripleSpec slit issue:

   Mark said that the TripleSpec slit motion has been increasingly balky for 
some time, and has gotten so unreliable of late that we have powered off 
the mechanism and fixed the slit in the 1.1 arcsecond position.  We need 
feedback from users on the impacts of this, both in the short and longer 
terms.  Suzanne said that we plan to leave the slit in this configuration 
until the summer, when we could open the dewar for repairs.  Alternatively, 
we may decide to leave the slit set to 1.1 arcseconds permanently.  In 
the past years, only about 15% of the users have moved the slit wheel.

   John Wilson and Matt Nelson described the situation in more detail, 
including the potential risks in restoring the slit wheel functionality. 
The device has a Geneva mechanism whose index switch is misbehaving and 
may have suffered a mechanical failure.  Whether it can be repaired will 
require warming the instrument and opening the dewar, both of which entail 
some risk of damage to the detector.

Discussion among the group led to the following action item:

===> ACTION:  Users and Users Committee members to provide Mark and Suzanne 
with their opinions and impacts of leaving the TripleSpec slit wheel 
permanently set to the 1.1 arcsecond position.


NIIS run next week (and no ToOs):

Bill said that he thinks that NIIS is good to go this Thursday night. 
It is also on the schedule again for next week.  Send target requests 
to Bill at  Suzanne added that because of instrument 
change logistics (which are hard when NIIS is on the floor), we will be 
suspending ToO requests from 1-8 March.


Upcoming instrument upgrades (DIS, echelle):

Suzanne listed the instrument upgrades, including the new echelle slitviewer 
camera (done), the new echelle grating (May), the new DIS B400 grating 
(March), and a new slitviewer camera for DIS (by next fall).  Users should 
check the instrument status webpage (see above) for current updates.


Summer shutdown schedule:

Mark announced that the summer shutdown will start during the July bright 
time, 23 July (a Monday), and conclude on 10 August (a Friday).  We will 
realuminize M2 and M3, and carry out normal annual maintenance on the
telescope, dome and instruments.



[open from previous meeting]:


[new actions from this meeting]:

===> ACTION:  Users and Users Committee members to provide Mark and Suzanne 
with their opinions about the impacts of leaving the TripleSpec slit wheel 
permanently set to the 1.1 arcsecond position.


Next meeting:  The next Users Committee phonecon will be on Monday 26 March 
at 8:30 AM Pacific Time.  The agenda and other materials will be sent 
to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1264 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to